# TestGenEval: A Real World Unit Test Generation and Test Completion Benchmark

### v1

@@ D.2.1 Test Generation Metrics
Pass@1: Pass@1 measures if any test in the generated test suite passes when run on the code under test. We include this to not penalize models that generate longer test suites.
@@
Mutation score: The main limitation of coverage and pass@1 is that they can be potentially gamed (a model can invoke all methods in the class under test without testing anything and achieve 100% coverage and 100% pass@1). To compute mutation score we inject synthetic bugs into the code under test. We choose which bugs we inject in accordance with best practices Derezinska & Halas (2014); Offutt et al. (1996). We then measure the percentage of bugs detected by the test suite (should pass on the original code and fail on the buggy code). Unlike other metrics, mutation score is much harder to game, however it is compute costly, as we have to execute the entire test suite for each bug. We rank by coverage, as models still have relatively low coverage across the board and it is more compute efficient for the community to run (we allow for TestGenEval to be run omitting mutation score).
@@ E.1.1 Test Generation
Model All@1 Pass@1 Cov Cov@Pass Mut Mut@Pass
@@ F.5.2 TestGenEvalLite
F.6 Qualitative Error Analysis
@@
Listing 10: Codestral first test method completion and code method
@@
Listing 11: Codestral last test method completion
@@
Listing 12: Codestral extra test completion (no asserts)
@@
Listing 13: Codestral first test completion (no asserts)
@@
Listing 14: Codestral extra test method completion
@@
Listing 15: Codestral first test method completion
@@
Listing 16: Codestral last test method completion
@@
Listing 17: Codestral first test method completion
@@
Listing 18: Codestral first test method completion
@@ F.7 Frontiers of TestGenEval
Listing 19: No solve problem (class with many nested methods)
@@
Listing 20: No solve problem (dealing with IO/Files)
@@
Listing 21: One solve problem (only GPT-4o)
@@
Listing 22: One solve problem (only GPT-4o)
@@
Listing 23: One solve problem (only Codestral 22B)

### v2

@@ D.2.1 Test Generation Metrics
Any pass@1: Any pass@1 measures if any test in the generated test suite passes when run on the code under test. We include this to not penalize models that generate longer test suites.
@@
Mutation score: The main limitation of coverage and pass@1 is that they can be potentially gamed (a model can invoke all functions in the file under test without testing anything and achieve 100% coverage and 100% pass@1). To compute mutation score we inject synthetic bugs into the code under test. We then measure the percentage of bugs detected by the test suite (should pass on the original code and fail on the buggy code). Unlike other metrics, mutation score is much harder to game, however it is compute costly, as we have to execute the entire test suite for each bug. We rank by coverage, as models still have relatively low coverage across the board and it is more compute efficient for the community to run (we allow for TestGenEval to be run omitting mutation score).
We use cosmic-ray [^5] https://github.com/sixty-north/cosmic-ray to generate mutants and use the default set of mutation operators [^6] https://github.com/sixty-north/cosmic-ray/tree/master/src/cosmic_ray/operators. This default set of operators follows best practices defined by the mutation testing community (Derezinska & Halas, 2014; Offutt et al., 1996). Cosmic ray has 565 stars and commonly in mutation testing research (Sánchez et al., 2022; Deb et al., 2024). We choose to not filter out mutation operators to achieve the most granular results possible (with no filtering there is only a 1.06% uncertainty in mutation score results).
@@ E.1.1 Test Generation
Model All@1 Any@1 Cov Cov@Pass Mut Mut@Pass
@@ F.5.2 TestGenEvalLite
F.6 Qualitative Model Comparison
We outline three cases where TestGenEval discriminates between GPT-4o, CodeStral and Llama 3.1 405B (only one model suceeds in each of these examples).
F.6.1 Example 1 - Test setup (only solved by GPT-4o)
class QuerySet(AltersData):
def __init__(self, model=None, query=None, ...):
def repr(self):
return "<%s %r>" % (self.__class__.__name__, data)
Listing 4: Query set class for managing data from database.
def test_queryset_repr(self):
queryset = QuerySet(model=None)
assertEqual(repr(queryset), "<QuerySet []>")
Listing 5: Cut GPT-4o generated test
from django.db import connection, models
def test_query_set_init(self):
model = models.Model()
qs = QuerySet(model=model)
self.assertEqual(qs.model, model)
Listing 6: Cut Llama 3.1 405B generated test
This is not the case with both Llama 3.1 405B and Codestral 22B. Both models ignore the empty case entirely and instead hallucinate invocations or imports of models in Django (instead these models should have mocked the model object and tested the null case). All models failing to properly mock the class under test’s dependencies lead to low coverage of the source file, despite GPT-4o generating passing tests all when other models fail.
F.6.2 Example 2 - Incorrectly mocking objects (only solved by Llama 3.1 405B)
class MigrationAutodetector:
def __init__(self, from_state, to_state, questioner=None):
self.existing_apps = {app for app, model in from_state.models}
Listing 7: Migration detection class to automatically detect changes needed between project states.
def test_changes():
class Model(models.Model):
pass
changes = migration.changes(graph={}, trim_to_apps=None)
self.assertEqual(len(changes), class="syntax-number">1)
Listing 8: Cut Llama 3.1 405B generated test
def test_init():
from_state = MagicMock(spec=ProjectState)
to_state = MagicMock(spec=ProjectState)
self.assertEqual(autodetector.from_state, self.from_state)
Listing 9: Cut GPT-4o generated test
F.6.3 Example 3 - Handling class dependencies (only solved by Codestral 22B)
class FieldGetDbPrepValueMixin:
class Exact(FieldGetDbPrepValueMixin, BuiltinLookup):
def process_rhs(self, compiler, connection):
Listing 10: Database classes to handle lookups and comparisons in queries.
def test_exact_lookup():
lhs = F(’field’)
lookup = Exact(lhs, ’value’)
self.assertIn(’%s’, ...)
Listing 11: Cut GPT-4o generated test
def test_exact():
field = Field()
lookup = Exact(field, ’value’)
self.assertEqual(lookup.lookup_name, ’exact’)
Listing 12: Cut Codestral 22B generated test
F.7 Qualitative Error Analysis
@@
Listing 13: Codestral first test method completion and code method
@@
Listing 14: Codestral last test method completion
@@
Listing 15: Codestral extra test completion (no asserts)
@@
Listing 16: Codestral first test completion (no asserts)
@@
Listing 17: Codestral extra test method completion
@@
Listing 18: Codestral first test method completion
@@
Listing 19: Codestral last test method completion
@@
Listing 20: Codestral first test method completion
@@
Listing 21: Codestral first test method completion
@@ F.8 Frontiers of TestGenEval
Listing 22: No solve problem (class with many nested methods)
@@
Listing 23: No solve problem (dealing with IO/Files)
@@
Listing 24: One solve problem (only GPT-4o)
@@
Listing 25: One solve problem (only GPT-4o)
@@
Listing 26: One solve problem (only Codestral 22B)